OpenStudy (pinksapphire):

Solve by using the best method... (x+2)(x+6)=x

OpenStudy (pinksapphire):

@tgawade

OpenStudy (pinksapphire):

@SolomonZelman

OpenStudy (pinksapphire):

@Hero

OpenStudy (anonymous):

first use FOIL method on left side , do you know FOIL method ?

OpenStudy (pinksapphire):

No...

OpenStudy (anonymous):

If (a + b)*(c+d) = a*c +a*d + b*c + b * d

OpenStudy (pinksapphire):

Should I subtract x?

OpenStudy (anonymous):

first so we will get (x+2) * (x +6) = x*x + x*6 + 2 *x+2*6 = x^2 + 6x + 2x + 12 = x2 + 8x + 12 right ?

OpenStudy (pinksapphire):

Okay...then what?

OpenStudy (anonymous):

So we have x^2 +8x + 12 = x Now to get rid of x from right hand side what we need to do ?

OpenStudy (anonymous):

Just subtract x from both sides & solve quadratic equation for x..

OpenStudy (anonymous):

So after subtracting x from both sides what do you get ?

OpenStudy (pinksapphire):

x^2+7x+12=0

OpenStudy (anonymous):

Correct !! , now do you know quadratic formula ?

OpenStudy (anonymous):

so what would be x = ?

OpenStudy (pinksapphire):

I hate the quadratic formula :(

OpenStudy (pinksapphire):

Give me a second...

OpenStudy (pinksapphire):

(-3,0) and (-4,0)